Find the third, sixth and eighth terms of the sequence A(n) = 6 + (n - 1)(-2). Step by step please, how do I get there ?
Answer by stanbon(75887)

Find the third, sixth and eighth terms of the sequence A(n) = 6 + (n - 1)(-2). Step by step please, how do I get there ?
---------------------
A(n) = 6 + (n - 1)(-2)
A(3) = 6 + 2(-2) = 2
A(6) = 6 + 5(-2) = -4
A(8) = 6 + 7(-2) = -8
--------------------------
Cheers,
Stan H.